# NeRFmentation: NeRF-based Augmentation for Monocular Depth Estimation

### v1

@@ 4.7 Ablation on NeRF-generated KITTI dataset
To demonstrate that NeRFmented MDE architectures generalize better on disturbed non-ideal test sets, we generate our own dataset by using novel views. More specifically, we train NeRFs on 111 sub-scenes from the KITTI Eigen test split. Then to filter out subpar scene representations, we set the LPIPS threshold to 0.3 and absolute relative error to 0.05, leaving 15 scenes for rendering. To render novel views, poses are disturbed randomly in the x and y directions and rotated randomly along each axis. This yields a test set with 1,218 dense RGB-D pairs.
In table 3 we see significant improvement in all metrics for all evaluated depth architectures using NeRFmented data. The trend of interpolation resulting in the best performance can be observed here as well.
[…]
$\uparrow$ $\uparrow$ $\uparrow$ $\downarrow$ $\downarrow$ $\downarrow$ $\downarrow$
AdaBins 0.761 0.961 0.988 0.183 0.987 4.248 0.211
+Reconstructed 0.901 0.985 0.995 0.128 0.645 3.424 0.143
+Interpolated 0.903 0.986 0.995 0.121 0.603 3.359 0.139
+Angled ±$\pm 3^{\circ}$ 0.892 0.983 0.994 0.132 0.682 3.541 0.147
DepthFormer 0.791 0.978 0.992 0.171 0.840 3.804 0.193
+Reconstructed 0.917 0.989 0.996 0.112 0.504 3.140 0.130
+Interpolated 0.922 0.989 0.996 0.106 0.457 3.009 0.126
 +Angled ±$\pm 3^{\circ}$ 0.917 0.989 0.996 0.112 0.512 3.098 0.129
BinsFormer 0.807 0.979 0.992 0.168 0.831 3.743 0.189
+Reconstructed 0.924 0.990 0.997 0.111 0.492 3.045 0.128
+Interpolated 0.925 0.990 0.996 0.105 0.459 2.994 0.125
+Angled ±$\pm 3^{\circ}$ 0.914 0.988 0.996 0.118 0.562 3.235 0.133
Table 3: Comparison of performances on the disturbed densified KITTI dataset. This table shows the performance of baseline and NeRFmented models evaluated on a dataset of images generated using NeRFs trained on scenes not part of the KITTI Eigen [9, 12] train split. Our method improves upon baseline performance independent of the architecture used. The best results are in bold dark green. The second best results are underlined light green.
@@ 5 Discussion
In this paper, we explore the idea of increasing the accuracy and robustness of monocular depth estimators by synthesizing new data from KITTI and introducing new views to train on. Our experiments show that for evaluation on sparse depth, we are able to achieve a performance close or better to the one when training on the vanilla dataset. For evaluation on disturbed densified depth we achieve better performance with all of the MDE networks we used as baselines. This shows that NeRFmented images are desirable if the sharpness of the depth maps is of interest. We also find that introducing new views greatly increases robustness since performance is much better when testing on an unseen dataset than without.
Limitations.
Our proposed method increases the accuracy and robustness of monocular depth estimators by remedying a lack of dense depth maps, limited viewing angles, and a general lack of scene diversity. However, given a dataset already rich with dense depth maps, varied viewing angles, and diverse scenes, NeRFmentation shows no improvement in comparison with baseline methods. The NYU-Depth V2 (NYU) dataset [31] is an example of this. It is an indoor dataset of various commercial and residential scenes captured using Microsoft Kinect. As such, the depth maps in the NYU dataset are dense. Furthermore, the dataset was recorded by manually panning the Kinect sensor. As a result, the scenes are captured from various viewpoints. We conjecture that the dense depth maps and varied poses constrain NeRFmentation’s potential for performance and robustness improvement. We provide experimental evidence for this in the supplementary material.
Future work.
NeRFs are good at compressing and interpolating data. However, they lack the ability to model regions that were unobserved during training. Thus, when we render novel views where parts of the image were unobserved during the training process, no constructive prediction can be made by the NeRF. As such these parts of the image are filled with noise, unnecessarily punishing the MDE. A masking strategy based on knowledge of observed areas of the scene is delivered in the supplementary material. Dynamic objects such as cars and people are incorrectly reconstructed due to the static world assumption made. Dynamic NeRF methods such as [20, 33, 19] could remedy this shortcoming, allowing for even higher fidelity data synthesis.
Acknowledgements
One of the authors is funded by 2023/00027/001 under the “Campaign financed by the European Union Next Generation - Recovery Plan - Competent Ministry - Spanish Government”.
@@ Acknowledgements
Appendix A Evaluation Metrics for MDE
@@
Appendix B Qualitative comparison with state-of-the-art methods on the KITTI and Waymo dataset
@@
Figures B.1 and B.2 provide additional qualitative results compared to the state-of-the-art methods AdaBins [6], Depthformer [16] and BinsFormer [15] on the KITTI [12] and Waymo [36] datasets. Qualitative improvements on KITTI are similar to those found on Waymo, our proposed method is able to predict increased details in objects such as trees and street signs, which can be especially noted in the top third part of the image.
Appendix C Qualitative comparison between the Reconstructed, Interpolated, and Angled
We compare the three proposed data augmentation strategies in conjunction with DepthFormer in Figure C.1. The differences between the different NeRFmented depth predictions on the Waymo dataset show that they all perform similarly. This suggests that the depth completion offered by NeRFs is the significant driver behind the performance gains.
Appendix D Scene mesh of NeRFs trained on KITTI
Using the built-in poisson mesh export in Nerfstudio [37] we generate a mesh of the NeRF scene. Figure D.1 shows a bird-eye’s view and a front-facing view of a 50-meter KITTI sub-scene. This 3D representation allows us to verify that the NeRF depth map predictions are sound and consistent.
Refer to caption Refer to caption
Bird’s-eye view Front-facing view
Figure D.1: Mesh generated with NeRF that was trained on KITTI [12] sub-scene. One can see that the underlying scene geometry is accurately learned. The bird’s-eye view shows what a single sub-scene looks like after the scene has been split into smaller 50-meter sub-scenes.
Appendix E Discussion: NeRFmentation on NYU-Depth V2
NYU-Depth V2 Dataset [31].
E.1 Training on NYU-Depth V2
The aforementioned MDE networks are trained on the original NYU-Depth V2 train split and NeRFmented versions of it using the approach described in Section 4.5. The relevant results are provided in Table E.1. NeRFmentation appears to hinder performance for both models and all novel view synthesis strategies. Unlike in the case of KITTI, for NYU-Depth V2, the introduction of synthetic images may not improve generalization to a more diverse set of poses since the original data is not particularly lacking in pose diversity. Instead, the introduction of noisy images that reduce the data quality becomes the main factor determining the net outcome of the augmentation. Additionally, our experiments on the KITTI dataset suggest that dense supervision coming from NeRFmentation is a large factor in why NeRFmentation improves on vanilla training. With NYU-Depth V2, the vanilla dataset already provides dense supervision. Therefore, this advantage of NeRFmentation is lost when using the NYU-Depth V2 dataset.
E.2 Zero-shot data transfer to Replica Dataset
To evaluate performance on a different indoor dataset, the Replica Dataset [34] is used. Replica consists of high-quality 3D indoor scene reconstructions, of which we choose the two validation scenes, and the one test scene from Bauer et al. [4] as our validation split. We modify the source code of the provided ReplicaRenderer and ReplicaViewer [34] to be able to render customized sequences of synthetic RGB-D images following a natural trajectory through each scene. NeRFmentation deteriorates generalization to the unseen data distribution of synthetic images from the Replica dataset, as seen in Table E.2 and Figure E.3. These results together with the NYU test set results show that the training data becomes too noisy for potential NeRFmentation benefits in generalization to unseen data distributions to be possible.
E.3 Ablation on NeRF-generated NYU-Depth V2 dataset
Similar to the procedure applied in the case of KITTI-based NeRFmented MDE architectures, we aim to demonstrate generalization to perturbed poses for the NYU-Depth V2 dataset. The same pose disturbance scheme as in KITTI is used, and the scheme is applied to NeRFs trained on the test scenes remaining after the same filtration process as in KITTI experiments to eliminate subpar scenes. To create a test set of approximately the same size as the original NYU-Depth V2 test set, up to 10 frames per scene are selected at equal intervals along the scene sequence, and new images are rendered using corresponding perturbed poses. With this approach, the perturbed test set consists of 700 images covering 73 scenes, capturing a diverse set of indoor environments. As demonstrated in Table E.3, the NeRFmented MDE models surpass the baseline AdaBins [6] models in the case of NeRFmentation using reconstructed views. In contrast, the DepthFormer [16] models do not exhibit a conclusive advantage for either the NeRFmented or the vanilla-trained models.
Augmentation $\delta_{1}$ $\delta_{2}$ $\delta_{3}$ REL RMS $\log_{10}$
$\uparrow$ $\uparrow$ $\uparrow$ $\downarrow$ $\downarrow$ $\downarrow$
AdaBins 0.903 0.984 0.997 0.103 0.364 0.044
+Reconstructed 0.877 0.977 0.994 0.118 0.414 0.049
+Interpolated 0.898 0.983 0.996 0.106 0.375 0.045
+Angled ±$\pm 3^{\circ}$ 0.884 0.981 0.996 0.112 0.387 0.047
DepthFormer 0.921 0.989 0.998 0.096 0.339 0.041
+Reconstructed 0.909 0.987 0.997 0.101 0.350 0.043
+Interpolated 0.909 0.987 0.997 0.101 0.353 0.043
+Angled ±$\pm 3^{\circ}$ 0.908 0.987 0.998 0.102 0.350 0.043
Table E.1: Comparison of performances on the NYU-Depth-v2 [31] dataset. This table shows the performance of AdaBins [6] and DepthFormer [16] models trained on the NYU-Depth V2 train split [9, 31] and compare it to the NeRFmented versions of the same models, using different novel view generation strategies. Both vanilla-trained and NeRFmented models were evaluated against the NYU-Depth V2 test split [9]. Our method yields reduced performance in most metrics for both models. The best results are in bold dark green. The second best results are underlined light green.
Augmentation $\delta_{1}$ $\delta_{2}$ $\delta_{3}$ REL RMS $\log_{10}$
$\uparrow$ $\uparrow$ $\uparrow$ $\downarrow$ $\downarrow$ $\downarrow$
AdaBins 0.512 0.921 0.982 0.271 0.821 0.103
+Reconstructed 0.214 0.576 0.895 0.526 1.572 0.179
+Interpolated 0.336 0.772 0.952 0.398 1.284 0.140
+Angled ±$\pm 3^{\circ}$ 0.300 0.713 0.916 0.418 1.297 0.157
DepthFormer 0.410 0.904 0.978 0.325 0.937 0.116
+Reconstructed 0.150 0.606 0.943 0.526 1.540 0.175
+Interpolated 0.115 0.542 0.908 0.582 1.634 0.190
+Angled ±$\pm 3^{\circ}$ 0.109 0.486 0.908 0.604 1.752 0.196
Table E.2: Comparison of performances on the Replica Dataset [34]. Models trained on the NYU-Test V2 train split [9, 31] are compared to the NeRFmented (Ours) versions of the same models evaluated against a subset of the Replica Dataset [34], a 3D scene reconstruction dataset. Our method impairs performance considerably on unseen synthetic indoor sequences, invalidating the possibility of improved generalization. The best results are in bold dark green. The second best results are underlined light green.
Augmentation $\delta_{1}$ $\delta_{2}$ $\delta_{3}$ REL RMS $\log_{10}$
$\uparrow$ $\uparrow$ $\uparrow$ $\downarrow$ $\downarrow$ $\downarrow$
AdaBins 0.706 0.907 0.960 0.794 0.750 0.094
+Reconstructed 0.732 0.912 0.964 0.622 0.691 0.083
+Interpolated 0.315 0.804 0.939 1.235 1.159 0.147
+Angled ±$\pm 3^{\circ}$ 0.311 0.804 0.936 1.347 1.129 0.148
DepthFormer 0.773 0.933 0.974 0.882 0.650 0.077
+Reconstructed 0.791 0.935 0.979 0.916 0.742 0.074
+Interpolated 0.797 0.934 0.978 0.948 0.731 0.072
+Angled ±$\pm 3^{\circ}$ 0.796 0.935 0.979 0.953 0.732 0.072
Table E.3: Comparison of performances on the perturbed NYU-Depth v2 [31] dataset. This table shows the performance of baseline and NeRFmented models evaluated on a dataset of images generated using NeRFs trained on scenes not part of the NYU-Depth v2 train split [31, 9]. Images are generated using random disturbations to original image poses. Our method does not yield an architecture-independent performance increase, rendering the robustness ablation for indoor MDEs inconclusive. The best results are in bold dark green. The second best results are underlined light green.
GT
Ours
GT
Ours
Figure E.1: Qualitative NeRF reconstruction results on the NYU-Depth V2 Dataset [31] We show original images from the NYU-Depth V2 dataset and the corresponding images generated using the trained and filtered NeRFs for the same camera poses. The reconstructed RGB images appear very similar to their real counterparts. The reconstructed depth maps are also very close to the ground truth depth maps. However, there is some noise apparent on the edges of objects in the reconstructed depth maps. Color scale: 0 (purple) to 10 meters (yellow).
RGB
Refer to caption Refer to caption Refer to caption Refer to caption
GT depth
Refer to caption Refer to caption Refer to caption Refer to caption
AdaBins [6]
Refer to caption Refer to caption Refer to caption Refer to caption
NeRFmented AdaBins [6]
Refer to caption Refer to caption Refer to caption Refer to caption
DepthFormer [16]
Refer to caption Refer to caption Refer to caption Refer to caption
NeRFmented DepthFormer [16]
Refer to caption Refer to caption Refer to caption Refer to caption
Figure E.2: Qualitative Results on the NYU-Depth V2 Dataset [31] We show the performance of NeRFmentation (Ours) compared to vanilla-trained AdaBins [6] and DepthFormer [16]. Both the NeRFmented models and vanilla-trained models have been trained on the NYU-Depth V2 train split and evaluated on the test split [9]. There is little discernible visual difference between the predictions of the vanilla-trained models and NeRFmented models. This is in line with the scores presented in Table E.1. Color scale: 1 (purple) to 6.5 meters (yellow).
RGB
Refer to caption Refer to caption Refer to caption Refer to caption
GT Depth
Refer to caption Refer to caption Refer to caption Refer to caption
AdaBins [6]
Refer to caption Refer to caption Refer to caption Refer to caption
NeRFmented AdaBins [6]
Refer to caption Refer to caption Refer to caption Refer to caption
DepthFormer [16]
Refer to caption Refer to caption Refer to caption Refer to caption
NeRFmented DepthFormer [16]
Refer to caption Refer to caption Refer to caption Refer to caption
Figure E.3: Qualitative Results on the Replica Dataset [35]. We show the performance of NeRFmentation (Ours) compared to vanilla-trained AdaBins [6] and DepthFormer [16]. Both the NeRFmented models and vanilla-trained models were trained on the NYU-Depth V2 train split [9, 31] and evaluated on a subset of the unseen Replica Dataset [35]. The NeRFmented models produce higher noise around edges of objects than the vanilla-trained models. This could be attributed to the noise introduced by NeRF-generated training data, as presented in Figure E.1. Color scale: 1 (purple) to 8 meters (yellow).
Appendix F Future Works: Masking
Due to the sparsity of depth maps of datasets that were recorded using a LiDAR sensor such as KITTI, NeRFs that are trained on these datasets have to inter- and extrapolate data that they have never seen before. Interpolating data for a NeRF is mostly unproblematic. However, NeRFs struggle at the scene borders of the training data as they have to extrapolate in these regions. Due to the fact that the level of supervision approaches zero in these regions, the NeRF learns to predict noise in these regions without any guarantee that the data there is useful. In the best case, the data corresponds to the actual real-world geometry, but in the worst case, it will predict values that could hurt the overall training procedure and thus need to be taken care of. This issue limits us from rendering more diverse poses for our NeRFmentation method, as any strong rotations or extreme translations will lead to noisy, but dense renders that we cannot mask out trivially without any extra steps.
In the following, we describe our proposed method to generate non-trivial binary occupancy masks that can be used to mask out unseen regions of NeRF-rendered depth maps during the training of a MDE network. These masks are designed to prevent the network from getting penalized for predicting incorrect values in pixels corresponding to these unseen regions.
Point cloud generation.
where
Mask generation.
When we generate an RGB-D image from a novel view, we store the pose used for each RGB-D image and split it into a rotation matrix $R_{W}^{NV_{k}}$ and a translation vector $t_{W}^{NV_{k}}$ to project the point cloud into the reference frame of the novel view $NV$ and using the intrinsic camera matrix $K_{C}$, we can then project the point-cloud into the image as a 2D projection and set the value for each pixel to 1 if it is occupied by data from at least one 3D point. This operation is the inverse of the point cloud generation process. It is described in the following equation 4:
Depth
Refer to caption Refer to caption
RGB
Refer to caption Refer to caption
Mask
Refer to caption Refer to caption
Masked Depth
Refer to caption Refer to caption
Masked RGB
Refer to caption Refer to caption
Figure F.1: NeRF reconstruction results with 15 degrees rotation and occupancy masks on KITTI [12]. We show novel views from KITTI that were generated by rotating 15 degrees to the left or right, starting from a training pose. We show the rendered, dense depth map and the rendered RGB image which clearly show artifacts in the border regions. Using the occupancy masks, it is possible to filter out these artifacts, but at the same time sparsify the depth maps.
$\text{mask}_{NV_{k}}(u,v)=\mathds{1}\{d_{NV_{k},uv}>0\},$
where
$d_{NV_{k},h}=K_{C}\cdot(R_{W}^{NV_{k}}\cdot\text{PC}_{s}+t_{W}^{NV_{k}}),$
$\text{PC}_{s}\in\mathbb{R}^{N\times 3}.$
For future work, we plan to investigate the impact of this proposed method. Our proposed method sparsifies each RGB-D image again and excludes the sky, a step that is already done during the evaluation on KITTI. As a result, we expect this method to not yield a significant improvement over our proposed vanilla NeRFmentation. However, we do expect substantial benefits for renders captured from extreme poses. Since these renders would deviate strongly from the data distribution of KITTI and would be too sparse for Waymo, we expect a notable performance loss on KITTI and no gain on Waymo. As a result, our extended method might not be beneficial for current datasets.
Figure F.1 shows rendered RGB-D images from challenging poses and how the occupancy masks can effectively remove these artifacts.
@@ Appendix F Future Works: Masking
Figure B.1: Qualitative Results on the KITTI [12] dataset. We show the performance of NeRFmentation (Ours Interpolated) compared to vanilla-trained AdaBins [6], DepthFormer [16] and BinsFormer [15] leading to improved predictions by using NeRF augmented data for introducing additional viewpoints in the training images. Both the NeRFmented models and vanilla-trained models have been trained on the KITTI Eigen [9, 12] train split and evaluated on test split. Color scale: 0 (purple) to 80 meters (yellow).

### v2

@@ 4.8 Ablation on NeRF-generated KITTI dataset
We believe our evaluation on KITTI and Waymo is limited by their sparse ground truth and limited number of viewing directions. To fully capture the improved generalization capabilities by NeRFmentation, we design our own synthetic test set with dense ground truth based on KITTI. We train NeRFs on 111 sub-scenes from the KITTI Eigen test split. Then, we filter out subpar scene representations by setting the LPIPS threshold to 0.3 and abs. rel. threshold to 0.05, leaving 15 scenes for rendering. Finally, we render novel views whose poses are randomly translated and rotated along each axis. This yields a test set with 1,218 dense RGB-D pairs.
In Tab. 4 we see significant improvement in all metrics for all evaluated depth architectures using NeRFmented data. This demonstrates that a non-ideal dense test set can better capture the value that is provided to the depth architectures by the NeRFmented data.
@@ 5 Discussion
In this paper, we explore the idea of improving the accuracy and generalizability of monocular depth estimators by synthesizing new data from existing data via neural radiance fields. Our evaluation shows that our method NeRFmentation greatly improved performance on the unseen Waymo and densified and perturbed KITTI dataset while retaining solid results on the base KITTI dataset. As classic augmentations hurt the performance when using NeRFmentation for out-of-distribution tasks, we argue that our presented method, NeRFmentation, should be considered a replacement for classic augmentations.
Limitations.  NeRFmentation improves monocular depth estimators by addressing sparse depth maps, limited viewing angles, and scene diversity issues. However, it shows no improvement for datasets already rich in these aspects, such as NYU-Depth V2 [28]. This indoor dataset, captured with Microsoft Kinect, provides dense depth maps from varied viewpoints. We hypothesize that these characteristics limit NeRFmentation’s potential for improvement (see supplementary material for evidence).
Future work.  NeRFs struggle with unobserved regions during training, producing noise in novel view renderings that unfairly penalize the MDE. A masking strategy for observed areas is provided in the supplementary material. Additionally, the static world assumption incorrectly reconstructs dynamic objects like cars and people. Dynamic NeRF methods [22, 30, 21] could potentially address this limitation, enabling higher fidelity data synthesis.
Acknowledgements.  One of the authors is funded by 2023/00027/001 under the “Campaign financed by the European Union Next Generation - Recovery Plan - Competent Ministry - Spanish Government”.
@@
  • $\diamond$
    Appendix 0.A provides details about the evaluation metrics of the monocular depth estimation algorithms.
  • $\diamond$
    Appendix 0.B shows an additional ablation about the nerfmented views saturation ablation on the KITTI and Waymo datasets.
  • $\diamond$
    Appendix 0.C gives additional qualitative results with state-of-the-art methods on the KITTI and Waymo dataset. Also, it compares NeRFmentation to DepthAnything.
  • $\diamond$
    Appendix 0.D introduces a qualitative comparison between the reconstructed, interpolated, and angled augmentation methods.
  • $\diamond$
    Appendix 0.E shows scene meshes of the NeRFs trained on the KITTI dataset.
  • $\diamond$
    Appendix 0.F presents a discussion about the proposed method applied to the NYU-Depth V2 dataset. The section discusses:
    • $\circ$
      Training on NYU-Depth V2.
    • $\circ$
      Zero-shot data transfer to Replica Dataset.
    • $\circ$
      Ablation on NeRF-generated NYU-Depth V2 dataset.
  • $\diamond$
    Finally, Appendix 0.G introduces the future works of NeRFmentation.
Appendix 0.A Evaluation Metrics for MDE
@@
Appendix 0.B NeRFmented views saturation ablation on the KITTI and Waymo dataset
We investigate the effect of adding increasing amounts of NeRFmented views to the base KITTI dataset when using the interpolation strategy with Depthformer in Tab. 0.B.1. We render views at 10 interpolated viewpoints between each pair of training poses. We then chose a random subset ranging between 1k to 50k NeRFmented images to add to the base 23k image KITTI dataset. Hyperparameters and iterations were kept constant for all runs. On the KITTI eigen test split, we see a small decrease in performance as more NeRFmented views are added. On the Waymo test set, we observe increasing performance as the added view count reaches 5-10k images and then a decrease in performance as the novel views become the majority of the training dataset. It is very interesting to note, that only 1k additional views (+4.3%) already lead to great gains on Waymo.
Table 0.B.1: Comparison of NeRFmented view count evaluated on KITTI and Waymo. We observe the effects of adding an increasing number of interpolated views to the base KITTI dataset. These results are for DepthFormer. Classic augmentation is used for 0k views and turned off for the others. We see a slight decrease in performance on KITTI as more views are added. The best results on Waymo are between 5k and 10k additional views. The best results are in bold dark green. The second best results are underlined light green.
# Extra Views $\delta_{1}\uparrow$ $\delta_{2}\uparrow$ $\delta_{3}\uparrow$ REL $\downarrow$ SQ. REL $\downarrow$ RMS $\downarrow$ RMSLOG $\downarrow$
KITTI [12]  • 0k 0.975 0.997 0.999 0.052 0.158 2.143 0.079
 • 1k 0.973 0.997 0.999 0.053 0.158 2.174 0.081
 • 5k 0.973 0.997 1.000 0.054 0.162 2.185 0.081
 • 10k 0.973 0.997 1.000 0.054 0.160 2.163 0.081
 • 20k 0.973 0.997 0.999 0.053 0.160 2.183 0.081
 • 50k 0.972 0.997 0.999 0.055 0.165 2.226 0.083
Waymo [32]  • 0k 0.545 0.885 0.958 0.217 1.949 7.911 0.304
 • 1k 0.690 0.922 0.974 0.177 1.522 6.976 0.251
 • 5k 0.718 0.933 0.978 0.172 1.441 6.570 0.238
 • 10k 0.733 0.931 0.977 0.167 1.443 6.683 0.234
 • 20k 0.722 0.928 0.976 0.171 1.473 6.764 0.240
 • 50k 0.701 0.918 0.972 0.176 1.578 7.041 0.250
@@ Appendix 0.B NeRFmented views saturation ablation on the KITTI and Waymo dataset
Appendix 0.C Qualitative comparison with state-of-the-art methods on the KITTI and Waymo dataset
@@
Fig. 0.C.1 and Fig. 0.C.2 provide additional qualitative results compared to the state-of-the-art methods AdaBins [5], Depthformer [17] and BinsFormer [18] on the KITTI [12] and Waymo [32] datasets. Qualitative improvements on KITTI are similar to those found on Waymo, our proposed method can predict increased details in objects such as trees and street signs, which can be especially noted in the top third part of the image.
@@
Figure 0.C.1: Qualitative predictions on the Waymo [32] dataset. We show the performance of NeRFmentation (Ours) compared to AdaBins [5], DepthFormer [17] and BinsFormer [18]. The color scale goes from 0 (purple) to 80 meters (yellow). Best viewed on a monitor zoomed in.
@@
Refer to caption Refer to caption Refer to caption Refer to caption
AdaBins
Refer to caption Refer to caption Refer to caption Refer to caption
NeRFm. AdaBins
Refer to caption Refer to caption Refer to caption Refer to caption
DepthF.
Refer to caption Refer to caption Refer to caption Refer to caption
NeRFm. DepthF.
Refer to caption Refer to caption Refer to caption Refer to caption
BinsF.
Refer to caption Refer to caption Refer to caption Refer to caption
NeRFm. BinsF.
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 0.C.2: Qualitative Results on the KITTI [12] dataset. We show the performance of NeRFmentation (Ours Interpolated) compared to vanilla-trained AdaBins [5], DepthFormer [17] and BinsFormer [18] leading to improved predictions by using NeRF augmented data for introducing additional viewpoints in the training images. Both the NeRFmented models and vanilla-trained models have been trained on the KITTI Eigen [8, 12] train split and evaluated on test split. Color scale: 0 (purple) to 80 meters (yellow).
0.C.1 Comparison with DepthAnything
DepthAnything is a promising series of novel foundation models for relative and metric depth estimation. For comparison, we choose DepthAnything v1 [37] over v2 [38], as weights for metric depth estimation fine-tuned on KITTI are only publicly available for v1. To evaluate NeRFmentation, we use DepthFormer trained on additional vertically translated views from KITTI but without classic augmentations.
Table 0.C.1: Evaluation of DepthAnything v1 and NeRFmentation on KITTI and Waymo datasets.
KITTI Waymo
REL $\downarrow$ $\delta_{1}\uparrow$ REL $\downarrow$ $\delta_{1}\uparrow$
DepthAnything v1 0.046 0.982 0.192 0.564
NeRFmentation 0.053 0.974 0.160 0.724
Figure 0.C.3: Zoomed visualizations of the predictions of DepthAnything v1 (left) and NeRFmentation (right) on the Waymo dataset.
As shown in Table 0.C.1, the performance of DepthAnything on in-distribution tasks (KITTI) is impressive. However, the motivation for NeRFmentation is to avoid dataset-specific overfitting. When testing on Waymo, especially the large difference in $\delta_{1}$ scores suggests that DepthAnything’s estimates are significantly worse more often than those of NeRFmentation. This is also supported by the finer details of NeRFmentation’s predictions shown in Figure 0.C.3. Therefore, we argue that NeRFmentation is more robust to unseen data than DepthAnything.
We also want to highlight that finetuning DepthAnything could benefit from NeRFmentation as well. Unfortunately, this experiment was not possible due to time constraints.
Appendix 0.D Qualitative comparison between the Reconstructed, Interpolated, and Angled augmentation strategies
We compare the three proposed data augmentation strategies in conjunction with DepthFormer in Fig. 0.D.1. The differences between the different NeRFmented depth predictions on the Waymo dataset show that they all perform similarly. This suggests that the depth completion offered by NeRFs is the significant driver behind the performance gains.
RGB DepthFormer NeRFm. DepthFormer (Reconstructed) NeRFm. DepthFormer (Interpolated) NeRFm. DepthFormer (Angled)
Figure 0.D.1: Qualitative DepthFormer [17] results on the Waymo [32] dataset, focusing on close-up details. Greatly improved detail compared to the baseline, but no apparent differences between reconstructed, interpolated, and angled data augmentation strategies. Color scale: 0 (purple) to 80 meters (yellow).
Appendix 0.E Scene mesh of NeRFs trained on KITTI
Refer to caption Refer to caption
Bird’s-eye view Front-facing view
Figure 0.E.1: Mesh generated with NeRF that was trained on KITTI [12] sub-scene. One can see that the underlying scene geometry is accurately learned. The bird’s-eye view shows what a single sub-scene looks like after the scene has been split into smaller 50-meter sub-scenes.
Using the built-in Poisson mesh export in Nerfstudio [33] we generate a mesh of the NeRF scene. Fig. 0.E.1 shows a bird’s-eye view and a front-facing view of a 50-meter KITTI sub-scene. This 3D representation allows us to verify that the NeRF depth map predictions are sound and consistent.
Appendix 0.F Discussion: NeRFmentation on NYU-Depth V2
NYU-Depth V2 Dataset [28].
0.F.1 Training on NYU-Depth V2
The aforementioned MDE networks are trained on the original NYU-Depth V2 train split and NeRFmented versions of it using the approach described in Section 4.5. The relevant results are provided in Tab. 0.F.1. NeRFmentation appears to hinder performance for both models and all novel view synthesis strategies. Unlike in the case of KITTI, for NYU-Depth V2, the introduction of synthetic images may not improve generalization to a more diverse set of poses since the original data is not particularly lacking in pose diversity. Instead, the introduction of noisy images that reduce the data quality becomes the main factor determining the net outcome of the augmentation. Additionally, our experiments on the KITTI dataset suggest that dense supervision coming from NeRFmentation is a large factor in why NeRFmentation improves on vanilla training. With NYU-Depth V2, the vanilla dataset already provides dense supervision. Therefore, this advantage of NeRFmentation is lost when using the NYU-Depth V2 dataset. Please refer to Fig. 0.F.1 and Fig. 0.F.2 for qualitative results.
GT
Ours
GT
Ours
Figure 0.F.1: Qualitative NeRF reconstruction results on the NYU-Depth V2 Dataset [28] We show original images from the NYU-Depth V2 dataset and the corresponding images generated using the trained and filtered NeRFs for the same camera poses. The reconstructed RGB images appear very similar to their real counterparts. The reconstructed depth maps are also very close to the ground truth depth maps. However, there is some noise apparent on the edges of objects in the reconstructed depth maps. Color scale: 0 (purple) to 10 meters (yellow).
0.F.2 Zero-shot data transfer to Replica Dataset
To evaluate performance on a different indoor dataset, the Replica Dataset [31] is used. Replica consists of high-quality 3D indoor scene reconstructions, of which we choose the two validation scenes, and the one test scene from Bauer et al. [3] as our validation split. We modify the source code of the provided ReplicaRenderer and ReplicaViewer [31] to be able to render customized sequences of synthetic RGB-D images following a natural trajectory through each scene. NeRFmentation deteriorates generalization to the unseen data distribution of synthetic images from the Replica dataset, as seen in Tab. 0.F.2 and Fig. 0.F.3. These results together with the NYU test set results show that the training data becomes too noisy for potential NeRFmentation benefits in generalization to unseen data distributions to be possible.
RGB
Refer to caption Refer to caption Refer to caption Refer to caption
GT Depth